OpenStudy (kayla2001):

51 kayla2001 y=6-x and y=x-2 a.) 2,4 b.) -4,2 c.) 4,2 d.) no solution

OpenStudy (anonymous):

Substitution.

OpenStudy (kayla2001):

I need this fast please

OpenStudy (anonymous):

Since it says y = x-2, then we can put that in the first equation.

OpenStudy (anonymous):

So it is x - 2 = 6 - x

OpenStudy (anonymous):

First, add x + x

OpenStudy (kayla2001):

I thought it was b or c

OpenStudy (anonymous):

That gives you 2x.

OpenStudy (anonymous):

Then add 2 + 6. @kayla2001

OpenStudy (kayla2001):

well am I right

OpenStudy (anonymous):

Just follow what I am doing. What is 2 + 6?

OpenStudy (kayla2001):

8

OpenStudy (holly00d1248):

I am agreeing so far

OpenStudy (anonymous):

Yes, now what's 8 \(\div\) 2?

OpenStudy (kayla2001):

4

OpenStudy (holly00d1248):

can i ask why you chose Substitution???

OpenStudy (anonymous):

Yes, so the value of x is 4.

OpenStudy (anonymous):

Because it is the easiest one to choose.

OpenStudy (anonymous):

*use not choose

OpenStudy (kayla2001):

its aaaaaaaaaaaaaaaaaaa

OpenStudy (anonymous):

@kayla2001 We can use the value of x in one of the equations to find the value of y.

OpenStudy (anonymous):

y=6-4 Can you figure this out?

OpenStudy (anonymous):

To find the value of y, you subtract 4 - 2, so what is the value of y? @kayla2001

OpenStudy (holly00d1248):

right coordinates wrong order Kayla try again

OpenStudy (kayla2001):

o c

OpenStudy (anonymous):

@kayla2001 Can you figure out 4 - 2 for the value of y? We already have 4 for the value of x.

OpenStudy (holly00d1248):

Whats o c

OpenStudy (kayla2001):

2

OpenStudy (holly00d1248):

so if X=4 and Y=2 whats the answer.

OpenStudy (holly00d1248):

@kayla2001 are you there

OpenStudy (anonymous):

Yes, what choice is your answer? @kayla2001